OpenStudy (anonymous):

Two supplementary angles have measures of 9x degrees and 3x degrees. What is the measure of the longer angle?

OpenStudy (danjs):

supplementary means they add up to a straight line or 180 degrees

OpenStudy (danjs):

3x + 9x = 180

OpenStudy (anonymous):

Thank you so much! very helpful!

OpenStudy (emmatassone):

\[x(3+9)=180\rightarrow x=15\]

OpenStudy (emmatassone):

longer angle= 15*9=135º